OpenStudy (anonymous):

help!! how do i solve for n in: (.9985atm)(.02L)= (0.082 atm*L/mole*K)(296K) sooo stumped!!!

OpenStudy (aaronq):

(.9985atm)(.02L)= n (0.082 atm*L/mole*K)(296K) n=[(0.082 atm*L/mole*K)(296K)]/[(.9985atm)(.02L)]

OpenStudy (anonymous):

i cant get a definite number/figure for n tho? i dont understand how to caclulate 0.082 atm*L/mole*K or what it means

OpenStudy (chmvijay):

its PV =nRT n=PV/RT = .9985atm *02L / 0.082 atm*L/mole*K *296K = find it atm*L/mole*K its just unit of gas constant don't worry about it

OpenStudy (chmvijay):

@missy101 got it ???:)

OpenStudy (chmvijay):

@aaaronq is wrong !!!

OpenStudy (aaronq):

my baddd i copied and pasted backwards

OpenStudy (chmvijay):

its all right it happens some times !!!:)

OpenStudy (anonymous):

when you go out on a limb, you sometimes fall off the tree. it's really minor anyway.

OpenStudy (toxicsugar22):

CAN YOU PLEAE HELP ME

OpenStudy (anonymous):

hmmm ok so theres no specific answer? i was told i needed to find one. . . dont see how its posible :/

OpenStudy (anonymous):

if you want the numbers, give me a sec.

OpenStudy (chmvijay):

8.2 10^-4 is answer LOL:)

OpenStudy (anonymous):

you want the number of gas particles in moles, or n. using the ideal gas law: n = P V / (R T) putting everything in atm, L, and K: n = (.9985 * .02) / (0.082 * 296) mol it comes out to be 8.2 x 10(-4) mol, just like chmvijay posted